# GenIR: Generative Visual Feedback for Mental Image Retrieval

### v1

@@ 1 Introduction
Recent Vision-language models (VLMs) have achieved decent results on standard text-to-image retrieval benchmarks wang2024qwen2 ; li2024llava . Despite this progress, transferring these capabilities into real-world applications remains challenging. One key limitation is that real human search behavior is often not one-shot or static; it unfolds through a sequence of actions, highlighting the necessity for interactive information retrieval (IIR) systems rui1999image ; white2009exploratory ; borlund2013interactive . Another limitation is that users frequently initiate a search to re-find the information they have seen before, which could be partial memory, vague clues, or vivid recall of the target images dumais2003stuff ; zhou2003relevance ; guo2018dialog . To address this scenario, we define Mental Image Retrieval (MIR)  [^2] The term is inspired by the Mental Image Reconstruction task koide2024mental and shares the definition of “mental image”, though their study purely from Neuroscience side which is different problem from ours., where users iteratively refine their queries based on mental image (i.e., an image in mind) to retrieve an intended image from an image database.
@@
Although MIR has not been explicitly formulated as a distinct task previously, as a subset of long-standing text-to-image IIR task thomee2012interactive ; guo2018dialog , some works have implicitly touched upon similar settings. ChatIR levy2023chatting positions a VLM (or ideally a human) as the active searcher. Seeing the initial query from the searcher, ChatIR uses Large Language Models (LLMs) to provide system feedback in question format to the user solely based on the textual dialogue history. The question is then answered by the searcher who has access to the ground-truth images (i.e., Mental Image). Next, the dialogue history will be appended with the question-answer pair and then be used as the search query. This implicitly positions ChatIR within the realm of MIR as a subset of interactive text-to-image retrieval where the human searcher holds the target image in memory. However, as shown in the bottom left of Figure 1, ChatIR has only verbal (text-based) feedback with no information from image space, resulting in generated question-answer pairs may be redundant or irrelevant to the query refinement. PlugIR lee2024interactive , as shown in the bottom right of Figure 1, advances this setup further by incorporating retrieval context—text captions of retrieved images into the query generation for subsequent rounds, aiming to produce more contextually relevant feedback and mitigate redundancy. Nevertheless, the major challenge exists, both methods remain constrained by significant limitations regarding feedback effectiveness yang2024tackling . Even across multiple interaction rounds, these methods rely heavily on indirect, verbal feedback derived solely from retrieval failures. Such feedback is often abstract and interpretability-poor, providing users with little actionable insight or potentially misleading clues for refining subsequent queries. In vision-language embedding spaces, such as CLIP radford2021learning , minor textual edits can cause unpredictable changes in retrieval outcomes, making query refinement inherently a trial-and-error process. Consequently, the feedback offered by existing conversational retrieval approaches inadequately expresses the AI system’s current understanding and fails to directly benefits users toward effective refinements. As ChatIR example shown in Figure 6, the verbal feedback for an image depicting a person wearing a motorcycle helmet: “a human is not wearing a hat”. Although literally true, such feedback fails to capture the visual salience of the helmet and may steer the user’s refinement toward irrelevant details, ultimately misleading the search process for images containing headwear.
@@
Beyond its utility at inference time, GenIR also supports dataset construction for studying MIR. Following the common practice of using VLM to play as the human searcher levy2023chatting ; lee2024interactive , we create an automated pipeline based on the GenIR framework. We present a multi-round dataset with each round consisting of a refined query, a generated synthetic image, and retrieved results with a correctness label. Our experiments demonstrate that GenIR outperforms exisitng MIR baselines, highlighting the significant advantage of using visual feedback over verbal feedback. Furthermore, our study indicates that our GenIR annotated query can result in better retrieval performance than annotation from ChatIR under the same retrieval setting. In summary, our contribution is as follows:
@@ 2.1 Chat-based Image Retrieval
Conversational image retrieval, which use chat-like feedback to improve query for text-to-image IIR has gained attention as a way to improve search performance mo2024survey . A foundational work, ChatIR levy2023chatting , demonstrated improved retrieval accuracy through multi-round chats, where an LLM poses questions answered by a human with target image access. ChatIR also contributed a multi-round chat dataset and highlighted the utility of multi-round interaction for retrieval tasks. However, both the performance of their method and the quality of the curated dataset were limited by feedback efficiency issue (redundancy or misleading) as we discussed in Section 1. PlugIR lee2024interactive advanced this idea by proposing a plug-and-play image captioning model to collect feedback from retrieved images. This yielded context-aware and non-redundant verbal feedback. However, both these two works and their related subsequent works yoon2024bi ; zhao2025chatsearch remain fundamentally limited in their capacity to share nuanced visual representation between user.
@@ 2.2 Generative Image for Image Retrieval
Diffusion models have achieved great success in image reconstruction, that is, given a target image, a text encoder is trained to output human-readable language or a latent representation as input to the diffusion model, aiming to generate an image close to target image wang2024visual ; wei2024diffusion . However, applying such models directly to MIR is non-trivial, as human users can hardly provide actual images based on their mental images. A more feasible attempt is Imagine-and-Seek li2024imagine , which involves a one-time process that uses an image captioning model to generate a text description from the target image and then feeds it into a text-to-image diffusion model to generate proxy image for retrieval. Yet, as discussed in the section 1, this single-round approach has been proven by multiple interactive retrieval works to be inferior in dealing with real-world applications guo2018dialog ; thomee2012interactive .
@@ 3.1 Task Formulation
We formally define the task of Mental Image Retrieval (MIR) as a subset of the text-to-image Interactive Information Retrieval thomee2012interactive . MIR inherits the nature of multi-round interaction from IIR, while only focusing on the case where the user has an internal mental image that can not be directly accessed by the retrieval system. From an Information Retrieval theory perspective, MIR does not consider the Exploratory Search where a searcher has never seen the searching target marchionini2006exploratory ; white2009exploratory , but focuses on Known-item Search where the searcher has seen and can recall or partially recall the target information wildemuth1995known ; ogilvie2003combining ; broder2002taxonomy . To define the task, we denote an image database $\mathcal{N}$, and let $I^{\text{target}}$ represent the image that the user has in mind. The retrieval task proceeds over multiple interaction rounds $t=1,2,...,T$. At each round, the user formulates a textual query $q_{t}$ intended to approximate the mental representation of $I^{\text{target}}$. Based on the query $q_{t}$, the retrieval system returns a candidate image $I^{\text{retrieved}}_{(t)}\in\mathcal{N}$. Additionally, the system provides feedback signals to the user, potentially benefiting subsequent refinements of the query to bridge discrepancies between the retrieved candidate image and the user’s mental image. The iterative process continues until the target image is successfully identified or a predefined maximum number of interaction rounds is reached.
@@ 4.3 Results and Analysis
Prediction feedback is not always better verbal
@@ 5 Dataset Contribution
As a byproduct of our experimental framework, we construct a multi-round dataset for MIR task. Algorithm 1 describes our automated dataset curation pipeline. Initially, the VLM formulates an initial textual description based on the target image. Subsequently, at each interaction round, the following steps are executed: (1) a synthetic image is generated via a diffusion-based text-to-image model conditioned on the current textual query (Line 6); (2) an image-to-image retrieval mechanism identifies the closest matching database image (Line 7); (3) correctness is automatically labeled based on retrieval outcome (Line 8); and (4) the VLM refines the textual query informed by visual discrepancies between the synthesized and target images(Line 10). Unlike ChatIR, our dataset centers on visual feedback: both the query issuer and the image retriever share their understanding through images. This shared visual grounding reduces redundancy and helps prevent the generation of misleading information.
@@ 6 Limitation
Our study has two primary limitations: First, our VLM simulation assumes users have a clear, fixed target image in mind, whereas real users often begin with only partial or fuzzy mental representations. Second, our framework doesn’t account for how mental images naturally evolve and clarify during the search process itself, as retrieval attempts often help users refine their own memory. Future work should include human studies that capture these dynamic aspects of memory retrieval to validate GenIR’s effectiveness in more naturalistic search scenarios. We leave more discussion on each point and future work to the Appendix E.
@@ D.4 Limitations and Future Work
D.5 Computational Cost vs. Performance Analysis
@@
D.5.1 Comparative Computational Analysis
@@
D.5.2 Optimizations and Efficiency Improvements
@@
D.5.3 Real-World Deployment Considerations

### v2

@@ 1 Introduction
Recent Vision-language models (VLMs) have achieved decent results on standard text-to-image retrieval benchmarks wang2024qwen2 ; li2024llava . Despite this progress, transferring these capabilities into real-world applications remains challenging. One key limitation is that real human search behavior is often not one-shot or static; it unfolds through a sequence of actions, highlighting the necessity for interactive information retrieval (IIR) systems rui1999image ; white2009exploratory ; borlund2013interactive ; yang2024rag . Another limitation is that users frequently initiate a search to re-find the information they have seen before, which could be partial memory, vague clues, or vivid recall of the target images dumais2003stuff ; zhou2003relevance ; guo2018dialog . To address this scenario, we define Mental Image Retrieval (MIR)  [^2] The term is inspired by the Mental Image Reconstruction task koide2024mental and shares the definition of “mental image”, though their study purely from Neuroscience side which is different problem from ours., where users iteratively refine their queries based on mental image (i.e., an image in mind) to retrieve an intended image from an image database.
@@
Although MIR has not been explicitly formulated as a distinct task previously, as a subset of long-standing text-to-image IIR task thomee2012interactive ; guo2018dialog , some works have implicitly touched upon similar settings. ChatIR levy2023chatting positions a VLM (or ideally a human) as the active searcher. Seeing the initial query from the searcher, ChatIR uses Large Language Models (LLMs) to provide system feedback in question format to the user solely based on the textual dialogue history. The question is then answered by the searcher who has access to the ground-truth images (i.e., Mental Image). Next, the dialogue history will be appended with the question-answer pair and then be used as the search query. This implicitly positions ChatIR within the realm of MIR as a subset of interactive text-to-image retrieval where the human searcher holds the target image in memory. However, as shown in the bottom left of Figure 1, ChatIR has only verbal (text-based) feedback with no information from image space, resulting in generated question-answer pairs that may be redundant or irrelevant to the query refinement. PlugIR lee2024interactive , as shown in the bottom right of Figure 1, advances this setup further by incorporating retrieval context—text captions of retrieved images into the query generation for subsequent rounds, aiming to produce more contextually relevant feedback and mitigate redundancy. Nevertheless, the major challenge exists, both methods remain constrained by significant limitations regarding feedback effectiveness. Even across multiple interaction rounds, these methods rely heavily on indirect, verbal feedback derived solely from retrieval failures. Such feedback is often abstract and interpretability-poor, providing users with little actionable insight or potentially misleading clues for refining subsequent queries. In vision-language embedding spaces, such as CLIP radford2021learning , minor textual edits can cause unpredictable changes in retrieval outcomes, making query refinement inherently a trial-and-error process. Consequently, the feedback offered by existing conversational retrieval approaches inadequately expresses the AI system’s current understanding and fails to directly benefit users toward effective refinements. As ChatIR example shown in Figure 6, the verbal feedback for an image depicting a person wearing a motorcycle helmet: “a human is not wearing a hat”. Although literally true, such feedback fails to capture the visual salience of the helmet and may steer the user’s refinement toward irrelevant details, ultimately misleading the search process for images containing headwear.
@@
Beyond its utility at inference time, GenIR also supports dataset construction for studying MIR. Following the common practice of using VLM to play as the human searcher levy2023chatting ; lee2024interactive , we create an automated pipeline based on the GenIR framework. We present a multi-round dataset with each round consisting of a refined query, a generated synthetic image, and retrieved results with a correctness label. Our experiments demonstrate that GenIR outperforms existing MIR baselines, highlighting the significant advantage of using visual feedback over verbal feedback. Furthermore, our study indicates that our GenIR annotated query can result in better retrieval performance than annotation from ChatIR under the same retrieval setting. In summary, our contribution is as follows:
@@ 2.1 Chat-based Image Retrieval
Conversational image retrieval, which use chat-like feedback to improve query for text-to-image IIR has gained attention as a way to improve search performance mo2024survey . A foundational work, ChatIR levy2023chatting , demonstrated improved retrieval accuracy through multi-round chats, where an LLM poses questions answered by a human with target image access. ChatIR also contributed a multi-round chat dataset and highlighted the utility of multi-round interaction for retrieval tasks. However, both the performance of their method and the quality of the curated dataset were limited by feedback efficiency issue (redundancy or misleading) as we discussed in Section 1. PlugIR lee2024interactive advanced this idea by proposing a plug-and-play image captioning model to collect feedback from retrieved images. This yielded context-aware and non-redundant verbal feedback yang2024tackling ; however, both of these works and their related subsequent works yoon2024bi ; zhao2025chatsearch remain fundamentally limited in their capacity to share nuanced visual representations with users.
@@ 2.2 Generative Image for Image Retrieval
Diffusion models have achieved great success in image reconstruction, that is, given a target image, a text encoder is trained to output human-readable language or a latent representation as input to the diffusion model, aiming to generate an image close to the target image wang2024visual ; wei2024diffusion ; ukkonen2020generating . However, applying such models directly to MIR is non-trivial, as human users can hardly provide actual images based on their mental images. A more feasible attempt is Imagine-and-Seek li2024imagine , which involves a one-time process that uses an image captioning model to generate a text description from the target image and then feeds it into a text-to-image diffusion model to generate a proxy image for retrieval. Yet, as discussed in the section 1, this single-round approach has been proven by multiple interactive retrieval works to be inferior in dealing with real-world applications guo2018dialog ; thomee2012interactive .
@@ 3.1 Task Formulation
We formally define the task of Mental Image Retrieval (MIR) as a subset of the text-to-image Interactive Information Retrieval thomee2012interactive . MIR inherits the nature of multi-round interaction from IIR, while only focusing on the case where the user has an internal mental image that can not be directly accessed by the retrieval system. From an Information Retrieval theory perspective, MIR does not consider the Exploratory Search where a searcher has never seen the searching target marchionini2006exploratory ; white2009exploratory , but focuses on Known-item Search where the searcher has seen and can recall or partially recall the target information wildemuth1995known ; ogilvie2003combining ; broder2002taxonomy . To define the task, we denote an image database $\mathcal{N}$, and let $I^{\text{target}}$ represent the image that the user has in mind. The retrieval task proceeds over multiple interaction rounds t=1,2,,Tt=1,2,...,T. At each round, the user formulates a textual query $q_{t}$ intended to approximate the mental representation of $I^{\text{target}}$. Based on the query $q_{t}$, the retrieval system returns a candidate image $I^{\text{retrieved}}_{(t)}\in\mathcal{N}$. Additionally, the system provides feedback signals to the user, potentially benefiting subsequent refinements of the query to bridge discrepancies between the retrieved candidate image and the user’s mental image. The iterative process continues until the target image is successfully identified or a predefined maximum number of interaction rounds is reached.
@@ 4.3 Results and Analysis
Prediction feedback is not always better than verbal
@@ 5 Dataset Contribution
As a byproduct of our experimental framework, we construct a multi-round dataset for MIR task. Algorithm 1 describes our automated curation pipeline where the VLM formulates an initial query from the target image, then at each round: (1) a synthetic image is generated (Line 6); (2) the closest database image is retrieved (Line 7); (3) correctness is labeled (Line 8); and (4) the VLM refines the query based on visual discrepancies (Line 10). Unlike ChatIR, our dataset centers on visual feedback where both parties share understanding through images, reducing redundancy and misleading information.
@@ 6 Limitation
Our study has two primary limitations: First, our VLM simulation assumes users have a clear, fixed target image in mind, whereas real users often begin with only partial or fuzzy mental representations. Second, our framework doesn’t account for how mental images naturally evolve and clarify during the search process itself, as retrieval attempts often help users refine their own memory. Additionally, while visual feedback is useful in the majority of cases, generated images can sometimes mislead refinement through hallucinations or detail misalignments (detailed failure mode analysis in Appendix D.5). Future work should include human studies that capture these dynamic aspects of memory retrieval to validate GenIR’s effectiveness in more naturalistic search scenarios. We leave more discussion on limitations and future work, including opportunities for RL-based system optimization, to Appendix E.
@@ D.4 Limitations and Future Work
D.5 Failure Mode Analysis of Visual Feedback
While our evaluation demonstrates that visual feedback is useful in 86% of cases (Section D), it is important to understand the failure modes where generated images can mislead the refinement process. We conducted a focused analysis of cases where GenIR failed to retrieve the target image, categorizing errors into three primary types based on how the generated visual feedback provided incorrect cues.
Table 3 presents our failure mode taxonomy with descriptions and representative examples from the MSCOCO dataset:
Table 3: Failure Mode Taxonomy: Categories of visual feedback errors that can mislead query refinement in GenIR
Error Type Description Example (MSCOCO)
Limited Improvement The generated image shows minimal visual change from the previous round despite query refinement, providing no new information to guide further refinement. Image 000000004952: Generated images from rounds 8 and 9 are nearly identical, stalling the refinement process.
Hallucination Content The generated image contains objects or scene elements that were not part of the user’s intent, diverting the refinement process toward irrelevant details. Image 000000020415: The 8th-round generated image hallucinates a shower head on the right side, an element entirely absent in the ground-truth, misguiding subsequent queries.
Retrieval-Detail Misalignment The generated image misrepresents details that are crucial for retrieval but may be perceived as trivial by observers, creating a gap between visual feedback and retrieval objectives. Image 000000026494: A long wooden bench in the ground-truth is incorrectly rendered as a single wooden chair. While seemingly minor, this detail is critical for accurate retrieval.
Analysis of Failure Patterns
Limited Improvement (stagnation) typically occurs in later rounds (rounds 7-10) when the model has exhausted its capacity to refine the generation based on incremental textual changes. This suggests a limitation in the diffusion model’s sensitivity to subtle query modifications.
Hallucination Content represents the most problematic failure mode, as it actively misleads users by introducing false visual elements. These hallucinations often stem from the diffusion model’s tendency to complete scenes with common co-occurring objects, even when not specified in the query.
Retrieval-Detail Misalignment highlights a fundamental challenge: what appears visually acceptable to humans may miss crucial details that distinguish the target in the retrieval space. This suggests the need for retrieval-aware image generation that prioritizes discriminative details.
Visualizations
Figure 8 illustrates representative examples of each failure mode with visual comparisons between ground-truth targets and problematic generated images:
Case 1: Limited Improvement (Image 04952)
Refer to caption Refer to caption
Round 8 Generated Round 9 Generated
Case 2: Hallucination Content (Image 20415)
Refer to caption Refer to caption
Ground Truth Round 8 Generated
Case 3: Retrieval-Detail Misalignment (Image 26494)
Refer to caption Refer to caption
Ground Truth (bench) Round 9 Generated (chair)
Figure 8: Visual examples of GenIR failure modes. Top: Stagnation between rounds 8-9. Middle: Hallucinated shower head. Bottom: Bench misrepresented as chair.
These visualizations demonstrate the current limitations of visual feedback generation and suggest future directions for improving feedback quality, such as retrieval-aware generation objectives or confidence-based hybrid feedback strategies.
D.6 Computational Cost vs. Performance Analysis
@@ D.5 Failure Mode Analysis of Visual Feedback
D.6.1 Comparative Computational Analysis
@@
D.6.2 Hybrid System: Balancing Performance and Efficiency
To address latency concerns while maximizing performance gains, we explored a hybrid approach that intelligently selects between fast Verbal Feedback and more powerful Visual Feedback. Our analysis shows that Visual Feedback uniquely succeeds in 22.3% of cases where Verbal Feedback fails, presenting an opportunity for strategic deployment.
Table 5 presents performance comparisons across different hybrid scenarios on MSCOCO:
  • Verbal Feedback: Baseline using text-only feedback (fastest, 2s/round)
  • Visual Feedback (Ours): GenIR with visual feedback (slower, 16s/round)
  • Hybrid Oracle: Perfect selection of when to use Visual vs. Verbal feedback
  • Random Select: Random selection using Visual Feedback for 22.3% of queries
Table 5: Hybrid System Performance: Comparison of Verbal Feedback, Visual Feedback, and Hybrid approaches (Hits@10%). The hybrid system uses Visual Feedback for 22.3% of queries and Verbal Feedback for 77.7%.
Dialog Verbal Visual Hybrid Random Verbal Verbal
[…]
0 74.48 89.71 93.40 77.88 +18.92 +3.40
1 82.73 93.11 95.97 85.06 +13.25 +2.33
2 85.83 95.00 97.43 87.89 +11.60 +2.06
3 87.97 95.97 97.91 89.76 +9.95 +1.79
4 89.13 96.51 98.20 90.81 +9.07 +1.68
5 89.96 96.85 98.30 91.50 +8.35 +1.54
6 90.49 97.14 98.59 91.98 +8.10 +1.49
7 90.98 97.48 98.64 92.42 +7.67 +1.44
8 91.27 97.67 98.79 92.69 +7.52 +1.42
9 91.80 97.72 98.79 93.13 +6.99 +1.33
10 92.33 98.01 98.88 93.62 +6.55 +1.29
Key findings from the hybrid system analysis:
  • Even random selection yields meaningful gains (+3.40% absolute improvement at Round 0) with only +3 seconds average latency increase per query on MSCOCO.
  • The oracle hybrid system demonstrates the upper bound of performance (+18.92% improvement at Round 0), showing significant room for improvement through intelligent query difficulty estimation.
  • In difficult cases where verbal feedback requires many rounds or fails to converge, GenIR’s ability to succeed in fewer rounds can actually be more time-efficient overall.
This analysis demonstrates that hybrid approaches offer a promising middle ground, providing substantial performance improvements with manageable latency overhead. Future work could develop learned policies to intelligently select between feedback modalities based on query characteristics.
D.6.3 Optimizations and Efficiency Improvements
@@
D.6.4 Real-World Deployment Considerations
@@ Appendix E Limitation and Future Work
Reinforcement Learning-Based System Optimization
While our current plug-and-play design intentionally provides flexibility for researchers to leverage state-of-the-art or domain-customized models, end-to-end optimization of the GenIR framework presents exciting opportunities for future work. We envision optimization from two complementary perspectives: the framework architecture and the unique dataset it produces.
GenIR as a Multi-Agent System. At its core, GenIR functions as a multi-agent system where the generator and retriever collaborate to solve retrieval tasks. By introducing explicit visual feedback into this loop, we create an interactive paradigm potentially amenable to established multi-agent optimization techniques. The modular design makes GenIR a flexible testbed where researchers can investigate how different agent capabilities impact system dynamics and optimization potential.
Trajectory-Rich Dataset for RL. Our dataset provides unique opportunities for system optimization through its rich, multi-step interaction trajectories:
  • Warm-up Training: Successful $(query\rightarrow generated\ image\rightarrow retrieval\ result)$ trajectories can provide initial supervised training before RL fine-tuning. This warm-up stage, standard practice for improving RL stability and sample efficiency, is particularly crucial for navigating the vast action space inherent to large generative models.
  • Trajectory-Based Optimization: Full interaction histories enable advanced trajectory-based optimization through process supervision (e.g., training reward models) rather than relying solely on delayed final rewards. This allows the system to learn from intermediate feedback quality, not just final retrieval success.
  • Diagnostic Analysis: Intermediate generated images provide unprecedented insight into how and why generative models misunderstand queries in task-oriented settings. This offers direct, actionable guidance for system-level debugging and task-specific optimization.
Optimization Strategy. We envision freezing the retriever while optimizing the image generator end-to-end. This design choice offers flexibility—the retriever need not be a trainable deep learning model, allowing use of fixed systems like Google Search or domain-specific rule-based search. For the generator, successful optimization depends on two objectives: (1) correctly reflecting text queries visually, and (2) distinguishing between retrieval targets and incorrect candidates in the image space, ensuring seamless component collaboration.
These outlined possibilities represent starting points for engaging the research community in multi-agent system optimization. We believe the community will discover many additional opportunities building on this foundation. Our work provides essential building blocks—a flexible framework and trajectory-rich dataset—to unlock the full potential of generative visual feedback in interactive AI tasks.